OpenStudy (anonymous):

line a intersects the y-axis at -1. its slope is -10/-9. plot another point. where would the other point be located?

OpenStudy (rhystic):

equation of a straight line is y = mx + b where m is the slope and b is the y-intercept. so you have y = 10/9 (x) + -1 (negative over negative makes positive so -10/-9 is same as 10/9) now just plug in a value for x (1 or 2 etc) and the nsolve the equation for y and you have a set of coordinates. for instance if you use x =1 then you have : 10/9*1+-1 = 1/9 so the coordinates of that point would be (1,1/9)